# Edge-SD-SR: Low Latency and Parameter Efficient On-device Super-Resolution with Stable Diffusion via Bidirectional Conditioning

### v1

@@ 1 Introduction
We are interested in the application of Stable Diffusion [16, 31] to Super Resolution (i.e. SD-SR), as it has been shown to produce steep improvements compared to previous SR approaches  [24, 15, 29]. Considering that SR is most widely used on mobile phones, where the vast majority of photos are taken nowadays, we focus on the very challenging task of creating a SD-SR model that can run on Edge devices (i.e. Edge SD-SR). Deploying such models on computationally limited devices, such as mobile phones, entails addressing several practical challenges due to the large model size and high latency. These issues are exacerbated for the SR task, as it operates on high-resolution and typically requires multiple model runs. For example, 4K×$\times$3K is a typical resolution on modern phones, while SD-SR typically generates $512\times 512$ patches per run, e.g. StableSR [24]. In this case, 88 model runs are required for tile-based generation, considering a $25\%$ overlap between tiles to avoid artifacts. Such a model runs in $\sim 900$ seconds on a high-end GPU (A100), making on-device deployment unfeasible. In this work, we aim to address both model size and latency while maintaining the model’s visual quality.
@@ 2 Related work
Diffusion for SR: we focus SR based on diffusion models [16], which have shown remarkable capabilities, albeit at the cost of high compute cost, which we address in this work. The impressive performance of SD for generation tasks resulted in early interest in the SR community. Early work [18] adopted SD for SR in some constrained settings like face and class-constrained SR. General formulations for in-the-wild images were later proposed, e.g. [17]. Often works leverage image priors of SD models for SR [14, 24]. For example, [24] exploited these priors by using adapters on top of a frozen SD model, achieving remarkable visual quality. More recent work [15] proposed scale distillation and produced state-of-the-art results with only one inference step. However, these methods build on top of SD1.5 [16], which is highly parameter and latency-inefficient for on-device deployment. We consider instead the orthogonal problem of training a compact model effectively for SD-SR. Notably, our method is compatible with scale distillation, and we leverage it to train a one-step model.
@@ 4 Experiments
Datasets Metrics LDM StableSR YONOS ResShift Edge-SD-SR (Ours)
DIV2K Valid RealESRGAN degradations FID $\downarrow$ 26.47 24.44 21.86 30.45 25.37
LPIPS $\downarrow$ 0.2510 0.3114 0.2310 0.3076 0.249
PSNR $\uparrow$ 23.32 23.26 24.74 24.62 24.10
SSIM $\uparrow$ 0.5762 0.5726 0.6428 0.6210 0.617
MUSIQ $\uparrow$ 62.27 65.92 70.30 63.58 69.58
RealSR LPIPS $\downarrow$ 0.3159 0.3002 0.2479 0.3279 0.278
MUSIQ $\uparrow$ 58.90 65.88 69.21 59.87 65.20
DRealSR LPIPS $\downarrow$ 0.3379 0.3284 0.2721 0.3870 0.292
MUSIQ $\uparrow$ 53.72 58.51 66.26 54.13 55.66
DPED-iphone MUSIQ $\uparrow$ 44.23 50.48 59.45 38.59 60.09
- # STEPS $\downarrow$ 200 200 1 4 1
- # GFLOPS $\downarrow$ 9820 47403 2042 2651 142
- # Params (M) $\downarrow$ 241 1063 960 173 169
@@ 4.2 Comparison with state-of-the art
In this section, we evaluate the performance of our proposed Edge-SD-SR model in the standard real image super-resolution setting targeting ×$\times 4$ scale factor and compare it to state-of-the-art. As our main point of contribution is diffusion-based SR models, we focus on comparisons to SD-based models in Table 1. It is important to note here that this comparison is not fair by design as we operate on a challenging setting with (1) a very small model and (2) only consider results obtained with one denoising step. In contrast, each of the compared methods only targets one of these aspects that we tackle collectively in our work. All SD-SR methods compared here consume $>2000$ GFLOPs, whereas our model requires $142$ GFLOPs. Specifically, the original LDM model [16] uses a similarly sized U-Net (but larger auto-encoder) and requires 200 denoising steps, yet we outperform this model in all datasets and metrics. StableSR [24] uses a much larger model and 200 denoising steps but focuses on encoder training and once again we outperform this model in $7$ out of $10$ comparison points. YONOS-SR [15] use a much larger model, but specifically optimize for inference speed to obtain a model running in only 1 denoising step. This is a very challenging comparison point given that a similar strategy is adopted in our case to tackle inference speed but we do that with a very limited computational budget and as such we fall behind YONOS-SR in several comparison points. More closely related to ours is ResShift [29], which also proposes a new conditioning strategy to rely more heavily on the LR image and uses only $4$ denoising steps. Notably, while we use a much smaller model size (considering the auto-encoder size as well), our proposed model outperforms ResShift in $8$ out of $10$ comparison points. In addition, we show quantitatively in Figure 2 that our model yields superior or on-par results while being significantly smaller and faster than state-of-the-art baselines. Collectively, these results support the effectiveness of the proposed approach to enable SD-SR models to readily run on-device.

### v2

@@ 1 Introduction
We are interested in the application of Stable Diffusion [17, 33] to Super Resolution (i.e. SD-SR), as it has been shown to produce steep improvements compared to previous SR approaches  [25, 16, 31]. Considering that SR is most widely used on mobile phones, where the vast majority of photos are taken nowadays, we focus on the very challenging task of creating a SD-SR model that can run on Edge devices (i.e. Edge SD-SR). Deploying such models on computationally limited devices, such as mobile phones, entails addressing several practical challenges due to the large model size and high latency. These issues are exacerbated for the SR task, as it operates on high-resolution and typically requires multiple model runs. For example, 4K×$\times$3K is a typical resolution on modern phones, while SD-SR typically generates $512\times 512$ patches per run, e.g. StableSR [25]. In this case, 88 model runs are required for tile-based generation. Specifically, to get a $4K\times 3K$ image we split the corresponding $1000\times 750$ LR image to $128\times 128$ patches with $25\%$ overlap between patches. This yields 88 patches that the model has to process sequentially. Such a model runs in $\sim 900$ seconds on a high-end GPU (A100), making on-device deployment unfeasible. In this work, we aim to address both model size and latency while maintaining the model’s visual quality.
@@ 2 Related work
Diffusion for SR: we focus SR based on diffusion models [17], which have shown remarkable capabilities, albeit at the cost of high compute cost, which we address in this work. The impressive performance of SD for generation tasks resulted in early interest in the SR community. Early work [19] adopted SD for SR in some constrained settings like face and class-constrained SR. General formulations for in-the-wild images were later proposed, e.g. [18]. Often works leverage image priors of SD models for SR [15, 25, 12, 30]. For example, [25] exploited these priors by using adapters on top of a frozen SD model, achieving remarkable visual quality. More recent work [16] proposed scale distillation and produced state-of-the-art results with only one inference step. However, these methods build on top of SD1.5 [17], which is highly parameter and latency-inefficient for on-device deployment. We consider instead the orthogonal problem of training a compact model effectively for SD-SR. Notably, our method is compatible with scale distillation, and we leverage it to train a one-step model.
@@ 4 Experiments
Datasets Metrics LDM StableSR YONOS DiffBir SeeSR ResShift Edge-SD-SR (Ours)
DIV2K Valid RealESRGAN degradations FID $\downarrow$ 26.47 24.44 21.86 40.42 31.93 30.45 25.37
LPIPS $\downarrow$ 0.2510 0.3114 0.2310 0.427 0.3843 0.3076 0.249
PSNR $\uparrow$ 23.32 23.26 24.74 20.94 21.19 24.62 24.10
SSIM $\uparrow$ 0.5762 0.5726 0.6428 0.493 0.5386 0.6210 0.617
MUSIQ $\uparrow$ 62.27 65.92 70.30 62.05 68.33 63.58 69.58
RealSR LPIPS $\downarrow$ 0.3159 0.3002 0.2479 0.3658 0.3009 0.3279 0.278
MUSIQ $\uparrow$ 58.90 65.88 69.21 64.85 69.77 59.87 65.20
DRealSR LPIPS $\downarrow$ 0.3379 0.3284 0.2721 0.4599 0.3189 0.3870 0.292
MUSIQ $\uparrow$ 53.72 58.51 66.26 61.19 64.93 54.13 55.66
DPED-iphone MUSIQ $\uparrow$ 44.23 50.48 59.45 - - 38.59 60.09
- # STEPS $\downarrow$ 200 200 1 50 50 4 1
- # GFLOPS $\downarrow$ 9820 47403 2042 $>2000$ $>2000$ 2651 142
- # Params (M) $\downarrow$ 241 1063 960 1717 2524 173 169
@@ 4.2 Comparison with state-of-the art
In this section, we evaluate the performance of our proposed Edge-SD-SR model in the standard real image super-resolution setting targeting ×$\times 4$ scale factor and compare it to state-of-the-art. As our main point of contribution is diffusion-based SR models, we focus on comparisons to SD-based models in Table 1. It is important to note here that this comparison is not fair by design as we operate on a challenging setting with (1) a very small model and (2) only consider results obtained with one denoising step. In contrast, each of the compared methods only targets one of these aspects that we tackle collectively in our work. All SD-SR methods compared here consume $>2000$ GFLOPs, whereas our model requires $142$ GFLOPs. Specifically, the original LDM model [17] uses a similarly sized U-Net (but larger auto-encoder) and requires 200 denoising steps, yet we outperform this model in all datasets and metrics. StableSR [25] uses a much larger model and 200 denoising steps but focuses on encoder training and once again we outperform this model in $7$ out of $10$ comparison points. YONOS-SR [16] use a much larger model, but specifically optimize for inference speed to obtain a model running in only 1 denoising step. This is a very challenging comparison point given that a similar strategy is adopted in our case to tackle inference speed but we do that with a very limited computational budget and as such we fall behind YONOS-SR in several comparison points. More closely related to ours is ResShift [31], which also proposes a new conditioning strategy to rely more heavily on the LR image and uses only $4$ denoising steps. Notably, while we use a much smaller model size (considering the auto-encoder size as well), our proposed model outperforms ResShift in $8$ out of $10$ comparison points. In addition, we show quantitatively in Figure 2 that our model yields superior or on-par results while being significantly smaller and faster than state-of-the-art baselines. Notably, we also outperform more recent SD-based SR work like DiffBir [12] and SeeSR [30]. Collectively, these results support the effectiveness of the proposed approach to enable SD-SR models to readily run on-device.